DIFFERENTIAL GEOMETRY/PDE SEMINAR

Monday, January 12, 2004 Padelford C-36 3:45 pm (Note: special day and time!) Existence of periodic water waves with vorticity

Walter Strauss (Brown University)

Consider classical inviscid water waves with non-trivial vorticity, with a free surface under the influence of gravity over a flat bottom. Then there exist global continua of 2D periodic traveling waves. The governing equations are Euler’s and the pressure is constant on the free surface. There is a global continuum of such solutions for every choice of speed c, spatial period L, relative mass flux p_{0} and vorticity function γ that satisfy a certain explicit inequality. The inequality is satisfied for any c, L and γ provided p_{0} is sufficiently small. The waves are symmetric around each crest, the period is the distance between successive crests, and the profiles are monotone between crest and trough. Each continuum extends from flat waves all the way to waves that have stagnation points. Thus there exist many rotational periodic traveling waves of large amplitude. These results have been obtained as joint work with Adrian Constantin. Almost all the previous work on this problem, notably by Stokes, Nekrasov, Levi-Civita, Keady, Norbury, Amick and Toland, has been restricted to irrotational flows.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.